# A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:

### v6

@@ 1.3 Governments’ tax setting choices
Before examining the effects of the GMT, we first solve the unconstrained optimisation problem in the absence of the GMT. The non-haven country $n$ maximises a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, whose owners reside exclusively in the non-haven. The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal value of public funds (MVPF) (Hebous and Keen, , 2023). For $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues (e.g., Janeba and Schjelderup, , 2023; Hindriks and Nishimura, , 2025).
@@
where $\Pi\equiv\int_{\underline{\pi}}^{\infty}\pi dF$ is the total true profits of MNEs.
Each of the havens $h$ only cares about its tax revenues (multiplied by the MVPF $\lambda$), as the affiliate located in $h$ is owned by their headquarters in the non-haven. This gives
@@
Two comments on (6) are in order. First, tax rates in all countries are rising in the profit-shifting costs $\delta$ and falling in the number of tax havens $H$. Second, they are rising in the MVPF, $\lambda$. A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the value of tax revenues vis-à-vis the post-tax profits of MNEs. The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well. [^15] Note the difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT. Johannesen, (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country. By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs; hence MNEs diversify their allocation of shifted profits across all havens.
@@ 2.1 Tax regimes
Table 1. Possible Regimes of the Non-Cooperative Tax Game.
@@
Table 1 shows possible combinations of tax policies of the non-haven and the havens. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In these regimes, if the non-haven adopts a single GMT rate, the havens set a uniform tax rate at the same level or higher and therefore earn no positive tax revenues. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven for small MNEs. Alternatively, if the havens set a single non-GMT or a single GMT rate, the non-haven never wants to split its tax rate, due to the constant tax base elasticity. Hence only five possible tax regimes remain.
Regime 0 is the unconstrained tax-competition equilibrium, where the GMT rate $t_{M}$ is non-binding in all countries, as discussed earlier. As $t_{M}$ rises, the equilibrium moves through four additional regimes, characterised as follows.
Proposition 1
(Equilibrium regimes with binding GMT)
Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal value of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$. As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterised by the following four regimes, with all symmetric havens choosing the same tax schedule.
@@
In Regime 1, when the GMT rate $t_{M}$ exceeds the havens’ unconstrained rate $t_{h}^{0}$, they must apply $t_{M}$ to large MNEs. The non-haven’s unconstrained rate remains above $t_{M}$, so it continues to set a single rate above $t_{M}$ for all MNEs. The havens then choose whether to apply $t_{M}$ also to small MNEs or use a split schedule. By committing to a single GMT rate, they can induce the non-haven to raise its uniform rate. This argument is strongest when the havens decide collectively, as assumed in our main model. [^16] Qualitatively, the same results hold when the havens make their first-stage decisions in a decentralised manner, as shown in Supplemental Appendix B. In this case, the parameter range of Regime 1 is smaller, since the commitment of a single haven induces the non-haven to raise its tax rate less than the simultaneous commitment of multiple havens. The havens trade off gains from inducing a higher non-haven rate against revenue losses from not undercutting the non-haven’s rate even more for small MNEs. In Regime 1, the first effect dominates, so they commit to the GMT rate for all MNEs, leading all countries to set higher uniform tax rates than in Regime 0.
@@
Notes: The figure shows the equilibrium tax rates in the non-haven in panel (a) and the haven in (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes in (c). While the levels of tax rates and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.
@@
Figure 1 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of the GMT rate $t_{M}$. The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b). The equilibrium regimes are summarised in panel (c). The solid (dotted) curves are the tax rates for large MNEs (small MNEs) in the tax-competition equilibrium under the GMT; the dashed horizontal lines are the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0). Tax rates in both countries generally rise as $t_{M}$ increases, but there are distinct patterns in the different regimes. At the switch from Regime 1 to 2, the havens begin to split their tax rates, and the rate for small MNEs falls discontinuously. At the switch from Regime 3 to 4, the non-haven splits its tax rate and discretely cut its tax rate for small MNEs to the level in the unconstrained Regime 0. The havens respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
2.2 Welfare effects of the GMT
We briefly summarise how the GMT introduction affects national and global welfare. Introducing a GMT rate $t_{M}$ slightly above the havens’ unconstrained one $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the havens. At the same time, profit shifting is reduced in comparison to the unconstrained Regime 0 due to a smaller tax differential. The sum of these effects must benefit the non-haven. It also benefits the havens, as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits.
Once $t_{M}$ exceeds a certain threshold within Regime 1, the relationship reverses for the havens, as the loss of shifted profits dominates. Introducing any GMT rate above this threshold benefits the non-haven but harms the havens, with the net effect raising global welfare. Supplemental Appendix C provides a formal proposition and proof of these results.
We also study the effects of marginal increases in the GMT tax rate $t_{M}$ and the GMT coverage rate $\phi$. An important result is that welfare in the non-haven and the world increases with $t_{M}$ in general, and the welfare gains become even larger as $\phi$ increases. A higher $\phi$ reduces the effectiveness of the split tax system and limits tax competition for small MNEs, thereby amplifying the welfare gains from raising $t_{M}$. See Supplemental Appendices D and E for formal discussions and Appendix F for graphical illustrations.
@@ 3 Quantitative implications
We now explore the quantitative importance of our theoretical results. We modify the model of the previous section by assuming that the havens decide tax policies in a decentralised manner (see footnote 16 and Supplemental Appendix B). The decentralised decision making leads to competition between the havens and thus makes it more likely that they split their tax rates. [^19] The new regime switching rate dividing Regimes 1 and 2 is smaller ($t_{M}^{b}<t_{M}^{1}$), and thus Regime 1 is less likely to emerge, as $t_{M}^{0},t_{M}^{b})\subset[t_{M}^{0},t_{M}^{1})$. This is because the commitment to a single GMT rate by a single haven does not induce the non-haven to raise its tax rate as much as the collective commitment by multiple havens, making gains from commitment smaller.
We calibrate the unconstrained model with no GMT (Regime 0) to match international profit-shifting data, as summarised in Table 2. The GMT coverage rate is set to $\phi=0.9$, following OECD, (2020) and our calculations from Orbis. [^20] See Supplemental Appendix A for details. The number of tax havens is set to $H=40$, as in Tørsløv et al., (2023). Combining their data with Country-by-Country Reporting (CbCR) yields total MNE pre-tax profits of $\Pi=4,623$ billion USD. The marginal value of public funds (MVPF), $\lambda$, and the cost parameter of profit shifting, $\delta$, are jointly calibrated to minimise the squared distance between the model-predicted non-haven tax rate and shifted profit share and their empirical counterparts. For non-targeted moments, our model slightly underestimates the havens’ tax rate, $t_{h}^{0}$, and the non-haven’s revenue losses, $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$, but overall fits the data reasonably well.
Table 2. Calibration of the Model.
@@
$\phi=0.9$ Profit share of MNEs covered by GMT OECD, (2020), Orbis
$H=40$ Number of tax havens Tørsløv et al., (2023)
$\Pi=4,623$bUSD Pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
$\lambda=2.12$ Marginal value of public funds Calibrated
$\delta=17.80$ Cost of profit shifting Calibrated
@@
Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al., (2023)
Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al., (2023), CbCR
@@
Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al., (2023)
@@
Notes: Data are from 2019. The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables’). The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world. There are $40$ haven countries across the world. The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates. Tørsløv et al., (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7). Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs. With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
@@
Using this calibrated version of our model, we examine the quantitative effects of GMT reforms, as summarised in Table 3. The key result is that a 15% GMT rate with partial coverage ($\phi=0.9$) yields the equilibrium in Regime 2 (Table 3(a)): the havens split their tax rates and set a lower tax rate ($10.1\%$) for small MNEs as compared to large MNEs ($15.0\%$) (Proposition 1(ii)). [^21] Above the regime switching rate of $t_{M}^{b}=0.094$, Regime 2 prevails. This result aligns with observations that low-tax countries like Ireland plan to retain their base rate (12.5% in Ireland’s case) while applying top-up taxes on large MNEs under the GMT (QDMTT).
Table 3. Quantitative Effects of GMT Reforms.
(a) Gains from a 15% GMT under $\phi=0.9$ Non-haven Haven World

### v7

@@ 2.3 Governments’ tax setting choices
Before examining the effects of the GMT, we first solve the unconstrained optimisation problem in the absence of the GMT. The non-haven country $n$ maximises a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, whose owners reside exclusively in the non-haven. The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal value of public funds (MVPF) (Hebous and Keen, 2023). For $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximising its tax revenues (e.g., Janeba and Schjelderup, 2023; Hindriks and Nishimura, 2025).
@@
where $\Pi\equiv\int_{\underline{\pi}}^{\infty}\pi dF$ are the total true profits of MNEs.
Each of the havens $h$ only cares about its tax revenues (multiplied by the MVPF $\lambda$), as the affiliate located in $h$ is owned by its headquarters in the non-haven. This gives
@@
Two comments on (6) are in order. First, tax rates in all countries are rising in the profit-shifting costs $\delta$ and falling in the number of tax havens $H$. Second, they are rising in the MVPF, $\lambda$. A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the value of tax revenues vis-à-vis the post-tax profits of MNEs. The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well. [^16] Note the difference between this result and the analysis in Johannesen (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT. Johannesen (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country. By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs; hence MNEs diversify their allocation of shifted profits across all havens.
@@ 3 Equilibrium tax regimes
Table 1
Possible regimes of the non-cooperative tax game.
@@
Table 1 shows possible combinations of tax policies in the non-haven and the havens. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In these regimes, if the non-haven adopts a single GMT rate, the havens set a uniform tax rate at the same level or higher and therefore earn no positive tax revenues. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven for small MNEs. Alternatively, if the havens set a single non-GMT or a single GMT rate, the non-haven never wants to split its tax rate, due to the constant tax base elasticity. Hence only five possible tax regimes remain.
Regime 0 is the unconstrained tax-competition equilibrium, where the GMT rate $t_{M}$ is non-binding in all countries, as discussed earlier. As $t_{M}$ rises, the equilibrium moves through four additional regimes, which are characterised as follows.
Proposition 1.
(Equilibrium regimes with binding GMT)  Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal value of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$. As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterised by the following four regimes, with all symmetric havens choosing the same tax schedule.
@@
In Regime 1, when the GMT rate $t_{M}$ exceeds the havens’ unconstrained rate $t_{h}^{0}$, they must apply $t_{M}$ to large MNEs. The non-haven’s unconstrained rate remains above $t_{M}$, so it continues to set a single rate above $t_{M}$ for all MNEs. The havens then choose whether to apply $t_{M}$ also to small MNEs or use a split schedule. By committing to a single GMT rate, they can induce the non-haven to raise its uniform rate. This argument is strongest when the havens decide collectively, as assumed in our main model. [^17] Qualitatively, the same results hold when the havens make their first-stage decisions in a decentralised manner, as shown in Section 5.2. In this case, the parameter range of Regime 1 is smaller, since the commitment of a single haven induces the non-haven to raise its tax rate less than the simultaneous commitment of multiple havens. The havens trade off gains from inducing a higher non-haven rate against revenue losses from not undercutting the non-haven’s rate even more for small MNEs. In Regime 1, the first effect dominates, so they commit to the GMT rate for all MNEs, leading all countries to set higher uniform tax rates than in Regime 0.
@@
Notes: The figure shows the equilibrium tax rates in the non-haven in panel (a) and the haven in (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes in (c). While the levels of tax rates and regime-switching rates vary with parameter values, the qualitative patterns remain unchanged.
@@
Fig. 2 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of the GMT rate $t_{M}$. The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b). The equilibrium regimes are summarised in panel (c). The solid (dotted) curves are the tax rates for large MNEs (small MNEs) in the tax-competition equilibrium under the GMT; the dashed horizontal lines are the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0). Tax rates in both countries generally rise as $t_{M}$ increases, but there are distinct patterns in the different regimes. At the switch from Regime 1 to 2, the havens begin to split their tax rates, and the rate for small MNEs falls discontinuously. At the switch from Regime 3 to 4, the non-haven splits its tax rate and discretely cut its tax rate for small MNEs to the level in the unconstrained Regime 0. The havens respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
4 Welfare effects of the GMT
Having fully described the tax-competition equilibrium in each regime, we can now turn to the welfare effects of introducing the GMT. As in Proposition 1, the qualitative nature of the propositions below holds independently of the marginal valuation of public funds, $\lambda$, and it includes revenue maximization ($\lambda\rightarrow\infty$) as a special case. Our discussion of the results will therefore focus on the implications that the GMT has on tax revenues.
We compare welfare levels in the tax-competition equilibrium to those in the unconstrained equilibrium, Regime 0. The results are summarized in:
Proposition 2.
(GMT introduction)  Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1, and assume the GMT coverage rate is not too low so that $\phi\in(3/4,1)$. Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
  • (i)
    a rise in welfare in the non-haven country for all Regimes 1 to 4.
  • (ii)
    a rise in welfare in the haven country for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
  • (iii)
    a rise in world welfare for all Regimes 1 to 4.
Proof: See B.2.
Introducing a small GMT slightly higher than the havens’ unconstrained rate $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the haven countries. At the same time, profit shifting is reduced in comparison to the no-GMT benchmark due to a smaller tax differential. The sum of these effects must certainly benefit the non-haven. It also benefits the havens, as long as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits. However, once the GMT rate exceeds a certain threshold, $t_{M}^{+}$, which still lies in Regime 1, the reduced tax differential and thus the fall in shifted profits will dominate from the perspective of the haven countries. Therefore, for all levels of the GMT above $t_{M}^{+}$, tax havens lose from the introduction of the GMT while the non-haven gains. [^20] This conflict of interest does not arise in (Johannesen, 2022, Proposition 6), where havens gain from the introduction of a GMT at any level, as long as profit shifting is not completely eliminated. In his setting where MNEs shift profits only to the lowest-tax country, all the havens end up setting their tax rate to zero. Therefore, revenues in the unconstrained tax competition equilibrium are zero for each haven, and any positive GMT rate generates a revenue increase for the representative haven. See also footnote 16.
In the remaining Regimes 2 to 4, the tax differential for large MNEs is always smaller than the one in Regime 0. As long as large MNEs covered by the GMT account for a sufficiently high share of aggregate profits, $\phi\in(3/4,1)$, the havens lose from the introduction of $t_{M}$ higher than $t_{M}^{+}$ due to the reduced profit shifting of large MNEs. Hence, the effects of introducing the GMT on tax revenues and welfare in the non-haven and the haven countries remain opposed to each other throughout Regimes 2 to 4.212121If small MNEs not covered by the GMT account for a sufficiently large share of aggregate profits so that $\phi\in(0,3/4)$, then the introduction of the GMT rate would instead benefit the havens in Regime 3. See B.2 for details. Finally, the gains to the non-haven exceed the losses (if any) to the havens, so that the GMT introduction raises tax revenues and welfare worldwide in all regimes. These aggregate welfare gains result from the higher tax rates in both countries and the reduced levels of profit shifting. Fig. C.1 in C illustrates the welfare effects summarised in Proposition 2.
After the GMT has been introduced, two further reform options exist: (i) a marginal increase in the GMT rate, $t_{M}$, and (ii) a marginal increase in the GMT coverage rate, $\phi$. An increase in $t_{M}$ generally increases welfare in the non-haven and the world, except at the regime-switching rates where either of the non-haven or the havens introduce a split tax system. For the non-haven country, the welfare effects of marginal increases in the GMT rate are non-monotonous. A higher $\phi$ has non-negative welfare effects for the non-haven and the world in all regimes, and it has non-positive welfare effects for the havens. Intuitively, a rise in $\phi$ limits tax competition for small MNEs, and therefore amplifies the welfare effects arising from the GMT. We fully study these reforms in D.1 and D.2 respectively.
@@ 5 Quantitative implications
5.1 Basic model
We calibrate the unconstrained model with no GMT (Regime 0) to match international profit-shifting data, as summarised in Table 2. The GMT coverage rate is set to $\phi=0.9$, following OECD (2020) and our calculations in Table 1. The number of tax havens is set to $H=40$, as in Tørsløv et al. (2023). Combining their data with Country-by-Country Reports (CbCR) yields total MNE pre-tax profits of $\Pi=4,623$ billion USD. The marginal value of public funds (MVPF), $\lambda$, and the cost parameter of profit shifting, $\delta$, are jointly calibrated to minimise the squared distance between the two model-predicted moments— the non-haven tax rate $t_{n}^{0}$ and the shifted profit share $\sum_{h}\theta_{h}^{0}$—and their empirical counterparts. For non-targeted moments, our model slightly underestimates the havens’ tax rate, $t_{h}^{0}$, and the non-haven’s revenue losses, $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$, but overall fits the data reasonably well.
Table 2
Calibration of the model.
@@ 5.1 Basic model
$\phi=0.9$ Profit share of MNEs covered by GMT OECD (2020), Orbis
$H=40$ Number of tax havens Tørsløv et al. (2023)
$\Pi=4,623$bUSD Pre-tax profits of MNEs Tørsløv et al. (2023), CbCR
$\lambda=2.1$ Marginal value of public funds Calibrated
$\delta=17.8$ Cost of profit shifting Calibrated
@@
Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al. (2023)
Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al. (2023), CbCR
@@
Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al. (2023)
Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al. (2023)
@@
Notes: Data are from 2019. The definitions of non-haven and haven countries follow Tørsløv et al. (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables’). The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world. There are $40$ haven countries across the world. The corporate tax rates for the representative the non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates. Tørsløv et al. (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7). Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs. With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
@@
Using this calibrated version of our model, we examine the quantitative effects of GMT reforms, as summarised in Table 3. The key result is that a 15% GMT rate with partial coverage ($\phi=0.9$) yields an equilibrium in Regime 1 (Table 3(a)), whereas a 16% GMT rate leads to an equilibrium in Regime 2 (Table 3(b)). At $t_{M}=0.16$, the havens split their tax rates and set a lower tax rate ($10.3\%$) for small MNEs as compared to large MNEs (16.0%) (Proposition 1(ii)). Indeed, the threshold GMT rate is $t_{M}^{1}=0.156$, above which Regime 2 prevails. As we will see below, the threshold level for the GMT to induce an equilibrium in Regime 2 is further reduced, if the havens set their tax rates in a decentralized way.
Table 3
Quantitative effects of GMT reforms.
(a) Gains from a 15% GMT with $\phi=0.9$ Non-haven Haven World
Tax rate after GMT introduction $20.5\%$ for all $15.0\%$ for all —–
(b) Gains from a 16% GMT with $\phi=0.9$ Non-haven Haven World
Tax rate after GMT introduction $20.6\%$ for all $\begin{cases}10.3\%&\text{for small}\\ 16.0\%&\text{for large}\end{cases}$ —–
Welfare change in bUSD (% change) $104.4$ ($4.2\%$) 8.8-8.8 ($9.9\%$) $95.6$ ($3.7\%$)
Revenue change in bUSD (% change) $165.0$ ($24.4\%$) 8.8-8.8 ($9.9\%$) $156.2$ ($20.3\%$)
Profit change in bUSD (% change) $422.3$ $(44.0\%)$ 422.3-422.3 $(-44.0\%)$ 0 $(0\%)$
[…]
Tax rate after GMT introduction $20.8\%$ for all $16.0\%$ for all —–
Welfare change in bUSD (% change) $114.5$ ($4.6\%$) 8.9-8.9 ($10.0\%$) $105.6$ ($4.1\%$)
Revenue change in bUSD (% change) $181.1$ ($26.8\%$) 8.9-8.9 ($10.0\%$) $172.2$ ($22.5\%$)
Profit change in bUSD (% change) $461.2$ $(48.1\%)$ 461.2-461.2 $(-48.1\%)$ 0 $(0\%)$
Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$); panel (b), from $t_{M}=0$ to $0.16$ with $\phi=0.9$; and panel (c), from $t_{M}=0$ to $0.16$ with full coverage ($\phi=1.0$). The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven. The regime-switching rate between Regimes 1 and 2 is $t_{M}^{1}=15.6\%$. Results are based on the calibrated parameter values described in Table 2. The welfare levels are normalised by the MVPF, $\lambda$. The ‘haven’ here refers to the sum of the $H=40$ haven countries.
For both $t_{M}=0.15$ and $0.16$, the non-haven’s revenues rise due to weaker tax competition (higher tax rates) and reduced profit shifting, while the havens’ revenues fall as the loss of shifted profits outweighs the gains from milder tax competition. The calibrated revenue gain in the non-haven (153–165 billion USD) is close to the combined gains for Europe and the U.S. in Baraké et al. (2022, Table 1) ($67+58=125$ billion EUR) and worldwide gains (149–156 billion USD) are in the same range as in Hugger et al. (2024, Section 8.2) (155–192 billion USD). Since falling shifted profits reduce MNEs’ private benefits, the non-haven’s welfare rises less than their revenues. [^22] In the havens without MNE headquarters, revenues coincide with welfare.
Do the quantitative effects of introducing GMT change if the coverage is extended? The effects of $t_{M}=0.15$ with full coverage ($\phi=1.0$) are unchanged since both the non-haven and the havens set a single tax rate for all MNEs, implying the same outcomes as in Table 3(a). By contrast, introducing $t_{M}=0.16$ with full coverage yields the non-splitting equilibrium in Regime 1 (Table 3(c)), rather than the splitting one in Regime 2 (Table 3(b)). By eliminating competition for small MNEs, the extended coverage increases revenue by about 16 billion USD and welfare by about 10 billion USD, both in the non-haven and worldwide. [^23] Comparing Tables 3(b) and 3(c), extending $\phi$ from 0.9 to 1.0 increases the non-haven’s revenues by 181.1165.0=16.1181.1-165.0=16.1 billion USD and their welfare by 114.5104.4=10.1114.5-104.4=10.1 billion USD. Corresponding numbers for the world are 172.2156.2=16.0172.2-156.2=16.0 billion USD and 105.695.6=10.0105.6-95.6=10.0 billion USD, respectively. These gains should be compared with the estimated compliance costs for small MNEs, however. Applying the results of Gaul et al. (2022) to our setting, the additional compliance costs are in the range of 10.5 billion USD. [^24] Using a survey of German MNEs (see Becker et al., 2012 for details on their characteristics), Gaul et al. (2022) estimate the average one-time expenses of implementing GMT per firm at 1.6 million EUR. Applying this to the number of small MNEs not covered by GMT in our database ($\approx$6,000; see Table A.1 in A), the total compliance costs of newly covered MNEs are 1.6 million EUR $\times$ 6,000 $\times$ 1.1 USD/EUR $\approx$ 10.5 billion USD. Subtracting this amount from the gains of a full-coverage GMT results in net revenue gains of 5.5 billion USD, but a small net welfare loss of –0.5 billion USD globally. [^25] Clifford et al. (2025) also estimate size-dependent compliance costs for German MNEs and similarly find that an increased coverage of the GMT would provide only limited gains for Germany.
5.2 Extensions
In this section we provide two extensions of our basic model and confirm the robustness of our quantitative findings.
5.2.1 Decentralised decision-making by tax havens
We modify the basic model by allowing the havens to decide tax policies in a decentralised manner. The decentralised decision-making leads to competition between the havens and thus makes it more likely that they split their tax rates. [^26] The new regime-switching rate dividing Regimes 1 and 2, denoted by $t_{M}^{b}$, is smaller than that under collective decision-making, i.e., $t_{M}^{b}<t_{M}^{1}$. Consequently, the parameter range in which Regime 1 emerges is narrower: $t_{M}^{0},t_{M}^{b})\subset[t_{M}^{0},t_{M}^{1})$. This is because the commitment to a single GMT rate by a single haven does not induce the non-haven to raise its tax rate as much as the collective commitment by multiple havens. This reduces the gains from committing to a single GMT rate for each haven, and increases the range of GMT rates that lead to a split tax equilibrium. A formal analysis of decentralized decision-making by tax havens is provided in E.1.
Using the calibrated parameter values in Table 2, we reassess the quantitative effects of the GMT reforms in Table 4.272727We can use the same calibrated parameter values as in the collective decision-making case because the analytical expressions for the targeted calibration moments, i.e., the non-haven’s rate $t_{n}^{0}$ and the share of shifted profits $\sum_{h}\theta_{h}^{0}$ in Regime 0, remain unchanged. Table 4(a) reports the results for a 15% GMT rate with partial coverage ($\phi=0.9$). In contrast to the collective decision-making case in Table 3(a), this reform yields an equilibrium in Regime 2: havens split their tax rates, setting a lower rate for small MNEs (10.1%) than for large MNEs (15.0%) (see Proposition 1(ii)). This result — a GMT rate of 15% leading the havens to adopt a split tax system — aligns with observations that low-tax countries like Ireland have announced to keep their base rate unchanged (in Ireland’s case at 12.5%) while applying additional taxes on large MNEs to comply with the GMT.
As a result of this response by the havens, the revenue and welfare gains from introducing a 15% GMT rate are smaller than under collective decision-making (Table 4(a) vs. Table 3(a)). Increasing the coverage rate from 90% to 100% at the 15% GMT rate eliminates the split tax system, as shown in Table 4(b), and brings revenue and welfare gains for both the non-haven and the world. [^28] Comparing Tables 4(a) and 4(b), extending $\phi$ from 0.9 to 1.0 increases the non-haven’s revenues by 153.0139.5=13.5153.0-139.5=13.5 billion USD and their welfare by 95.687.2=8.495.6-87.2=8.4 billion USD. Corresponding revenue and welfare gains for the world are 149.3135.7=13.6149.3-135.7=13.6 billion USD and 92.083.4=8.692.0-83.4=8.6 billion USD, respectively. These numbers are comparable to those under collective decision-making (see the last paragraph of Section 5.1).
Table 4
GMT reforms under decentralised decision-making by tax havens.
(a) Gains from a 15% GMT with $\phi=0.9$ Non-haven Haven World